OpenStudy (anonymous):

Write the equation of a line with slope m = 3 and including point (3, 5) in slope-intercept form.

OpenStudy (ajprincess):

Let the equation of the line be y=mx+c To find c plug in the values y=5, x=3 and m=3 in y=mx+c. Can u do it @flyinglikeabird??

OpenStudy (anonymous):

c is -1?

OpenStudy (anonymous):

err -4

OpenStudy (anonymous):

3x3 = 9 + -4 = 5

OpenStudy (ajprincess):

yup c=-4. So ur equation becomes y=3x-4. Getting it?

OpenStudy (anonymous):

I never really understand algebra. It all seems like made up crap. I see what you did, but what is c?

OpenStudy (ajprincess):

c is y-intercept

OpenStudy (anonymous):

Y=mx+b 5=3(3)+b 5=9+b B=4

OpenStudy (ajprincess):

@ninab731 b=-4 5=9+b 5-9=b -4=b

OpenStudy (anonymous):

Ahh, okay. It all makes sense how you did it, don't understand why it is that way...

OpenStudy (ajprincess):

y=mx+c is the general slope-intercept form

OpenStudy (anonymous):

I get it, just feels like I am missing something. But thanks for your help!

OpenStudy (ajprincess):

yw:) if anythng is nt clear plz ask